OpenStudy (anonymous):

I will give you a medal if you help me!!!! The equation y = 281(1.01)x is a model for the population of the United States y, in millions of people, x years after the year 2000. Estimate when the United States population will reach 400 million people

OpenStudy (anonymous):

putting y=400 you get x = 400/(281*1.01) = 1.4 i.e 1.4 yrs after 2000 i.e. sometime in 2001.... to be precise on 4/10*365 = 146th day of 2001 that would be 26th may 2001

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

:) happy to help